OpenStudy (lgbasallote):

this seems tricky =_= Prove: \[\frac{2 \tan x}{1 + \tan^2 x} = \sin (2x)\]

OpenStudy (lgbasallote):

i tried changing 1 + tan^2x to sec^2 x but it did nothing :/

OpenStudy (ajprincess):

change tanx to sinx/cosx and try

OpenStudy (lgbasallote):

i dont see it o.O

OpenStudy (lgbasallote):

\[\frac{\frac{2\sin x}{\cos x}}{1 + \frac{\sin^2 x}{\cos^2 x}}\]

myininaya (myininaya):

Clear the compound fraction action! :)

OpenStudy (lgbasallote):

lol didnt recognize you :p why change? haha and compound frawhat?

OpenStudy (lgbasallote):

srsly...i thought a new mod was in town

myininaya (myininaya):

Multiply top and bottom by cos^2(x)

OpenStudy (ajprincess):